# Beyond the Surface: A Global-Scale Analysis of Visual Stereotypes in Text-to-Image Generation

### 12 Jan 2024 v1

@@ 4 Analyzing Stereotypical Depictions in Generated Images
4.1 Do generated images reflect known stereotypes?
Prior research has shown that generative models have a tendency to exacerbate the bias against minority groups Bianchi et al. (2023b). We uncover commonly observed patterns in the generated images across different identity groups at a large scale by grounding our evaluations in existing textual stereotype resources like SeeGULL. We use the identified visual stereotypes for evaluating and identifying whether already known stereotypical attributes are visually represented in the images of different identity groups. We only consider identity groups that have more than one associated visual stereotypes for our analysis.
We find the likelihood of a stereotypical attribute $attr_{s}$ being present in any image of the identity group $id$, $\mathbb{L}(attr_{s},id)$, as follows. If a stereotypical attribute $attr_{s}$ was shown $n$ times per identity group across all 15 images and selected $k$ times by the annotators, then $\mathbb{L}(attr_{s},id)=\frac{k}{n}$. Figure 3 presents examples of the most likely stereotypes associated with certain identity groups as annotated by different annotators. We observe that the attributes with the highest likelihood for an identity group, also align with the known stereotypes present in SeeGULL. For example, the attributes ‘dark’, ‘thin’, ‘skinny’, and ‘underweight’, the most likely attributes depicting Sudanese individuals, are also known stereotypes in SeeGULL. The attribute ‘poor’ which has a likelihood measure of 0.5 for Bangladeshi, i.e., approximately 50% of images representing Bangladeshis contained a representation of ’poor’, is also an annotated stereotype in SeeGULL. Similarly, attributes ‘beaner’, ‘brown’, ‘festive, ‘dark skin’, and ‘sombrero’ are stereotypes in SeeGULL for Mexicans, and are also present in the images representing a Mexican person. An Indian person is often represented as ‘dark, ‘brown’, and ‘religious’ which are its associated visual stereotypes in SeeGULL.
4.2 Are some identity groups depicted more stereotypically than others?
Although we observe that images of identity groups contain known stereotypical attributes, it is also crucial to understand if representations of some identity groups tend to be more stereotypical than others. To analyze this better, we compute the ‘stereotype tendency’ $\theta_{id}$ for each identity group. We measure the mean likelihood of a stereotype being present in any image representing an identity group $\mathbb{L}({stereo},id)$ and compare it with the likelihood of a randomly selected non-stereotypical attribute being depicted visually in the same set of images $\mathbb{L}({random},id)$.
@@ 4.2 Are some identity groups depicted more stereotypically than others?
gives us the overall ‘stereotypical tendency’ of the generated images representing an identity group. Higher the ratio for an identity group, greater is the likelihood of a stereotypical representation.
@@
We observe that on average, the visual representation of any identity group is thrice as likely to be stereotypical than non-stereotypical, i.e., the visual stereotypical attributes associated with an identity group are thrice as likely to appear in their visual representation when compared to randomly selected visual non-stereotypical attributes. We also compute $\theta_{id}$ for all 135 identity groups. Table 1 presents the $\theta_{id}$ and the likelihood of images being stereotypical over non-stereotypical for 30 different identity groups. More specifically, images representing Togolese, Zimbabwean, Malian, Guyanese, Iraqi, etc., only contained stereotypical attributes when compared to random attributes. Images representing Nigerians and Sudanese were 27 times more likely to contain stereotypical attributes than randomly selected non-stereotypes. Representations of Ghanaians, Nepalese, and Chinese individuals have an approximately 5 to 6 times higher likelihood of containing stereotypical attributes. We will publicly release the dataset with all the annotated attributes.
Identity $\mathbb{L}$(stereo, id) $\mathbb{L}$(random, id) $\theta_{id}$
Togolese 0.350 0 N/A
Zimbabwean 0.322 0 N/A
Malian 0.289 0 N/A
Guyanese 0.155 0 N/A
Sierra Leonean 0.153 0 N/A
Guatemalan 0.139 0 N/A
Kosovar 0.116 0 N/A
Iraqi 0.114 0 N/A
Swedes 0.100 0 N/A
Danish 0.095 0 N/A
South Sudanese 0.091 0 N/A
Gabonese 0.052 0 N/A
Mauritanian 0.033 0 N/A
Greeks 0.030 0 N/A
Kuwaiti 0.025 0 N/A
Jordanian 0.016 0 N/A
Bhutan 0.016 0 N/A
Moroccan 0.008 0 N/A
Ecuadorian 0.006 0 N/A
Thailand 0.005 0 N/A
Liberian 0.333 0.004 75.00
Panamanian 0.241 0.004 54.37
Lebanese 0.190 0.004 51.53
Mauritian 0.175 0.005 36.83
Sudanese 0.379 0.014 27.35
Nigerian 0.075 0.003 27.29
Libyan 0.243 0.009 25.53
Egyptian 0.127 0.005 24.82
Laotian 0.191 0.009 21.56
Table 1: Likelihood of the representation of an identity group being stereotypical based on the ‘stereotypical tendency’ $\theta_{id}$ for the default representation of the identity group $(id)$.
4.3 Is it feasible to automatically detect stereotypes in images?
Annotating stereotypes in images at scale can often times be resource-intensive and time-consuming. Therefore, we also identify the feasibility of using automated methods employing captioning models for stereotype detection. We compare the results with and without using visual stereotypes as a reference. Without visual components to ground the stereotypes, the automated techniques detect a lot of culturally subjective attributes like ‘attractive’, ‘smart’, etc., which are not inherently visual but stereotypically associated with identity groups. However, on using visual attributes as a reference, our approach uncovers only the ‘visual’ stereotypes for an identity group. These stereotypes also have a high likelihood $L(attr_{s},id)$ of being present in the images as marked by the annotators 44.69% of the time. Figure 3 shows the most salient visual stereotypes associated with the identity group. Attributes like ‘sombrero’, ‘dark’, and ‘brown’ were the most salient visual stereotypes for Mexicans; ‘poor’ was highly salient with Bangladeshi, ‘dark’ and ‘thin’ for Sudanese, and ‘elegant’ for French. These attributes were also marked as being present by the annotators. However, our approach also identifies other stereotypical attributes which are not necessarily depicted in the images, e.g., attributes like ‘cow’, and ‘elephant’ for Indians. This could either be a limitation of our approach of detecting stereotypes in the generated captions or could be due to existing errors/biases in the generated captions themselves. Further analysis is required to tease out the errors and biases propagated by different components of such a system.
5 Offensiveness in the Depiction of Identity Groups
@@
Visual representation of some stereotypes can be more offensive than others. For example, the stereotypical depiction of an identity group as ‘poor’ can be considered more offensive when compared to the visual depiction of a stereotype ‘rich’ for another identity group. Therefore, it is also important to understand the offensiveness of these images, and whether certain identity groups have a more offensive representation when compared to others.
@@
The SeeGULL dataset has offensiveness rating for each stereotypical attribute associated with an identity group. We make use of this offensiveness score and infer an overall offensiveness score $O(id)$ for representation of identity groups. Let ${O}(attr_{s},id)$ denote the offensiveness score of a stereotypical attribute $attr_{s}$ associated with an identity group $id$ in SeeGULL. The mean offensiveness score for each identity group is then
@@
Figure 4: Offensiveness of the generated images across different countries. The depth of the color increases with the offensive nature of the images.
6 Stereotypical Pull in Generated Images: a Case Study
Figure 5: ‘Stereotypical Pull’: For minority groups, the generative models have a tendency to ‘pull’ the generation of images towards an already known stereotype even when prompted otherwise. The red lines indicate ‘stereotypical’ attributes; the blue lines indicates ‘non-stereotypical attributes’. The numbers indicate the mean cosine similarity score between sets of the image embeddings.
@@ 6 Stereotypical Pull in Generated Images: a Case Study
Prior research has shown that Text-to-Image models can have a ‘very homogenized lens’ when representing certain identities Qadri et al. (2023). We conduct a deeper analysis to better understand this across different identity groups and their generated images. If T2I models consistently generate images that align with the stereotypical representations of an identity group, despite being given neutral or non-stereotypical prompts, we term this phenomenon as ‘stereotypical pull.’ This points to the tendency of T2I models to revert to stereotypical depictions, reflecting inherent biases within the models. We illustrate the stereotypical pull across various identity groups using both quantitative and qualitative analysis.
We prompt the T2I models to generate images for 135 identity groups across three different sets by using the following prompts:
  • Default Representation (d): A/An $id$ person; where $id$ denotes the identity group.
@@
  • Stereotypical Representation (s): (i) A/An $id$ person described as $attr_{s}$, (ii) A photo/potrait of a/an $id$ person. described as $attr_{s}$; where $attr_{s}$ is the visual stereotypical attribute associated with the identity group $id$ in SeeGULL.
@@
  • Non-Stereotypical Representation (ns): (i) A/An $id$ person described as $attr_{ns}$, (ii) A photo/potrait of a/an $id$ person described as $attr_{ns}$; where $attr_{ns}$ is a visual attribute not associated with $id$ in SeeGULL.
We generate 15 images per prompt.
Analysing Diversity of Images Globally.
We compute the mean pairwise cosine similarity between the CLIP embedding representations across the above sets of images for identity groups. We use the following three measures.
@@
[…]
Tables  23 and  4 present the mean cosine similarity between the sets of images for all 135 identity groups across all stereotyped and non-stereotyped attributes. We observe that the default representation of any identity group has a higher similarity score with the ‘stereotyped’ images compared to the ‘non-stereotyped’ images for 121 out of 135 identity groups indicating an overall ‘pull’ towards generating stereotypical looking images, irrespective of the identity group. Moreover, for identity groups from global south, the mean similarity score ‘Mean Sim’ across all three measures is constantly higher indicating an overall lack of diversity in their representations even when prompted otherwise.
Disparate Image Diversity for Different Nationalities.
Figure 5 visually demonstrates the ‘stereotypical pull’ for the generated images of the identity groups ‘Bangladeshi’, ‘Swiss’, ‘Ethiopian’, and ‘Australian’ across two dimensions: (i) wealth, and (ii) physical attributes. The default representation of ‘a Bangladeshi person’, looks very similar to the sets of images of ‘a poor Bangladeshi person’ (stereotype), as well as ‘a rich Bangladeshi person’ (non-stereotype). However, the visual representation of ‘a Swiss person’ is quite distinct from the visual representation of ‘a poor Swiss person’ (non-stereotype) and ‘a rich Swiss person’ (stereotype). We observe this pattern even for physical characteristics, where images of ‘an Ethiopian person’, ‘a skinny Ethiopian person (stereotype)’, and ‘a rich Ethiopian person’ (non-stereotype) are all similar looking when compared to that of ‘an Australian person’, ‘a skinny Australian person’ (non-stereotype), and ‘an overweight Australian person’ (stereotype) which are visually quite distinct. The generative models have a tendency to ‘pull’ the image generation towards stereotypes for certain identities even when prompted otherwise.
[…]
Identity Group S(d, s) S(d, ns) S(s, ns) Mean Sim
South Sudanese 0.9824 0.964 0.9609 0.9691
Nigerien 0.9807 0.961 0.9608 0.9675
North Korean 0.9702 0.968 0.956 0.9648
Djiboutian 0.9719 0.9627 0.9596 0.9647
Myanmar 0.9720 0.9603 0.9551 0.9624
Bolivian 0.9688 0.9598 0.9587 0.9624
Mauritanian 0.9736 0.9575 0.9555 0.9622
Malawian 0.9837 0.9514 0.9491 0.9614
Bhutanese 0.9786 0.9546 0.9494 0.9609
Mongolian 0.9719 0.9554 0.9543 0.9605
Table 2: ‘Stereotypical Pull’ observed across generated images of sample identity groups. The default representations of images are more similar to their ‘stereotypical’ representations. Moreover, identity groups from global south consistently have an overall higher mean similarity score S($\cdot$) across their default (d), stereotypical (S), and non-stereotypical (ns) attributes indicating a ‘pull’ towards generating stereotypical images.
@@ 7 Conclusion
In this work, we present a comprehensive framework for evaluating Text-to-Image models with a focus on ‘regional stereotypes.’ Leveraging existing stereotype benchmarks in textual resources, we perform a three-fold evaluation of generated images. Firstly, we distinguish between inherently visual stereotypes, which can be represented in an image, and culturally subjective stereotypes. Subsequently, we identify visual stereotypes prevalent in the generated images from T2I models across 135 identity groups by conducting a large-scale annotation task. The annotated images, featuring highlighted visual markers of stereotypes, will be made publicly available for further research. Additionally, we quantify the offensiveness of the generated images across different identity groups, and investigate the feasibility of automatically detecting stereotypes in images. Through an extensive case study, we also demonstrate that generated images for historically marginalized groups exhibit a more visually ‘stereotypical’ appearance, even when Text-to-Image models are explicitly prompted with non-stereotypical attributes. We hope that the presented framework and the dataset will provide valuable insights for understanding and addressing regional stereotypes in visual content generation.
@@
In this work, we evaluate the stereotypes generated in visual depictions of people, based on their identity associated with a geographical location. We acknowledge that this is a complex notion of identity and will have overlap with other facets of a person’s identity, such as race, which is not a focus of this work. We also emphasize that this work does not cover all possible visual depictions of stereotypes and thus, it’s methods or dataset should not be used as a standard to deem model generations free of stereotypes. This work is an initial step towards a systematic, global-scale evaluations of T2I models for stereotypical depictions, and we hope for future work to build on this to attain more coverage of potential stereotypical and harmful depictions in generations.
@@ Acknowledgements
We thank Utsav Prabhu, and Kathy Meier-Hellstern for their feedback, and advice.

### 22 Feb 2024 v2

@@ 4 Study 1: Stereotypical Depictions
4.1 Stereotypes Identified through Human Annotations
Do generated images reflect known stereotypes?
We use the identified visual stereotypes for evaluating whether already known stereotypical attributes are visually represented in the images of different identity groups. We only consider identity groups that have more than one associated visual stereotypes for our analysis. We find the likelihood $\mathbb{L}(attr_{s},id)$ of a stereotypical attribute $attr_{s}$ being present in any image of the identity group $id$ as follows. If a stereotypical attribute $attr_{s}$ was shown $n$ times per identity group across all 15 images and selected $k$ times by the annotators, then $\mathbb{L}(attr_{s},id)=\frac{k}{n}$. Figure 3 presents examples of the most likely stereotypes associated with certain identity groups as annotated by different annotators. We observe that the attributes with the highest likelihood for an identity group, also align with the known stereotypes present in SeeGULL. For example, ‘dark’, ‘thin’, ‘skinny’, and ‘underweight’, the most likely attributes depicting Sudanese individuals, are also known stereotypes in SeeGULL. The attribute ‘poor’ which has a likelihood measure of 0.5 for Bangladeshi, i.e., approximately 50% of images representing Bangladeshis contained a representation of ’poor’, is also an annotated stereotype in SeeGULL. Similarly, attributes ‘beaner’, ‘brown’, ‘festive, ‘dark skin’, and ‘sombrero’ are stereotypes in SeeGULL for Mexicans, and are also present in the images representing a Mexican person. An Indian person is often represented as ‘dark, ‘brown’, and ‘religious’ which are its associated visual stereotypes in SeeGULL. We note that the terms ‘brown’, ‘black’ etc., in this work, as identified in images do not imply race, and instead are about appearance, or appearance-based observed race Hanna et al. (2020); Schumann et al. (2023), or even the colors themselves.
Are some identity groups depicted more stereotypically than others?
It is crucial to understand if representations of some identity groups tend to be more stereotypical than others. To analyze this better, we compute the ‘stereotypical tendency’ $\theta_{id}$ for each identity group. We measure the mean likelihood of a stereotype being present in any image representing an identity group $\mathbb{L}({stereo},id)$ and compare it with the likelihood of a randomly selected non-stereotypical attribute being depicted visually in the same set of images $\mathbb{L}({random},id)$.
@@ 4.1 Stereotypes Identified through Human Annotations
gives the ‘stereotypical tendency’ of the generated images. Higher the ratio, greater is the likelihood of its stereotypical representation.
@@
We observe that on average, the visual representation of any identity group is thrice as likely to be stereotypical than non-stereotypical, i.e., the visual stereotypical attributes associated with an identity group are thrice as likely to appear in their visual representation when compared to randomly selected visual non-stereotypical attributes. We also compute $\theta_{id}$ for all 135 identity groups. We observe that images representing Togolese, Zimbabwean, Swedes, Danish, etc., only contained stereotypical attributes when compared to random attributes, i.e., $\theta_{id}$ is infinite or N/A; whereas images representing Nigerians were 27 times more likely to contain stereotypical attributes than randomly selected non-stereotypes. (Please refer to the Tables 1 and 2 in the Appendix for all scores).
Figure 4: ‘Stereotypical Pull’: The generative models have a tendency to ‘pull’ the generation of images towards an already known stereotype even when prompted otherwise. The red lines indicate ‘stereotypical’ attributes; the blue lines indicates ‘non-stereotypical attributes’. The numbers indicate the mean cosine similarity score between sets of image embeddings.
How offensive are the depictions of different identity groups?
@@
Visual representation of some stereotypes can be more offensive than others. For example, the stereotypical depiction of an identity group as ‘poor’ can be considered more offensive when compared to the visual depiction of a stereotype ‘rich’. We investigate the offensiveness of images and whether certain identity groups have a more offensive representation compared to others.
@@
Using the offensiveness rating of stereotypical attributes in SeeGULL, we infer an overall offensiveness score $O(id)$ for representation of identity groups. Let ${O}(attr_{s},id)$ denote the offensiveness score of a stereotypical attribute $attr_{s}$ associated with an identity group $id$ in SeeGULL. The mean offensiveness score for each identity group is then
@@
4.2 Stereotypes Identified through Automated Methods
We check the feasibility of using automated methods employing captioning models for stereotype detection, and compare the results with and without using visual stereotypes as a reference. Without visual stereotypes to ground the evaluations, the automated techniques detect non-visual attributes like ‘attractive’, ‘smart’, etc., for identity groups. However, using visual attributes as a reference, our approach uncovers more objectively visual stereotypes for identity groups. These stereotypes also have a high likelihood $L(attr_{s},id)$ of being present in the images as marked by the annotators 44.69% of the time. Figure 3 shows the most salient visual stereotypes associated with the identity group. Attributes like ‘sombrero’, ‘dark’, and ‘brown’ were the most salient visual stereotypes for Mexicans; ‘poor’ was highly salient with Bangladeshi, ‘dark’ and ‘thin’ for Sudanese, and ‘elegant’ for French. These attributes were also marked as being present by the annotators. This approach also identified stereotypical attributes which were not necessarily depicted in the images, e.g., attributes like ‘cow’, ‘elephant’ for Indians. This could be a limitation in our automated approach or existing errors/biases in the generated captions themselves. Further analysis is required to tease out error and biases propagated by different components of such a system.
5 Study 2: Stereotypical Pull
Figure 5: ‘Stereotypical Pull’ observed across different identity groups. Y-axis is the similarity $S(\cdot)$ between stereotyped (s) and non-stereotyped (ns) images ($S\text{(}s,ns))$. X-axis represents the difference in the deviations of the stereotypical (s) and the non-stereotypical (ns) images from the default (d) representations $(S\text{(}d,s)-S\text{(}d,ns))$.
@@ 5 Study 2: Stereotypical Pull
Prior research has shown that Text-to-Image models can have a ‘very homogenized lens’ when representing certain identities Bianchi et al. (2023); Qadri et al. (2023). We conduct a deeper analysis to better understand this across the generated images of different identity groups. We define ‘stereotypical pull’ for any identity group as a Text-to-Image model’s inclination to generate images aligning with the stereotypical representations of an identity group when presented with (i) neutral prompts, and (ii) explicit non-stereotypical prompts. This points to the model’s tendency to revert to stereotypical depictions, reflecting its inherent biases.
We use the below sets of prompts, to generate 15 images per prompt for 135 identity groups and demonstrate the prevalence of stereotypical pull.
  • Default Representation (d): ‘A/An $id$ person; where $id$ denotes the identity group’.
@@
  • Stereotypical Representation (s): (i)‘A/An $id$ person described as $attr_{s}$’, (ii) ‘A photo/potrait of a/an $id$ $attr_{s}$ person’; where $attr_{s}$ is the visual stereotypical attribute associated with the identity group $id$ in SeeGULL.
@@
  • Non-Stereotypical Representation (ns): (i) ‘A/An $id$ person described as $attr_{ns}$’, (ii) ‘A photo/potrait of a/an $id$ $attr_{ns}$ person’; where $attr_{ns}$ is a visual attribute not associated with $id$ in SeeGULL.
Figure 4 visually demonstrates the ‘stereotypical pull’ for the generated images of the identity groups ‘Bangladeshi’, ‘Swiss’, ‘Ethiopian’, and ‘Australian’. For demonstrative purposes, we select opposing stereotypical and non-stereotypical attributes. The default representation of ‘a Bangladeshi person’, looks very similar to the sets of images of ‘a poor Bangladeshi person’ (stereotype), as well as ‘a rich Bangladeshi person’ (non-stereotype). However, the visual representation of ‘a Swiss person’ is quite distinct from the visual representation of ‘a poor Swiss person’ (non-stereotype) and ‘a rich Swiss person’ (stereotype). We observe this pattern even for physical characteristics, where images of ‘an Ethiopian person’, ‘a skinny Ethiopian person (stereotype)’, and ‘a rich Ethiopian person’ (non-stereotype) are all similar looking when compared to that of ‘an Australian person’, ‘a skinny Australian person’ (non-stereotype), and ‘an overweight Australian person’ (stereotype) which are visually quite distinct. The generative models have a tendency to ‘pull’ the image generation towards stereotypes for certain identities when prompted with neutral prompts, and non-stereotypical prompts.
@@ 6 Conclusion
In this work, we present a comprehensive framework for evaluating Text-to-Image models with a focus on ‘regional stereotypes.’ Leveraging existing stereotype benchmarks in textual resources, we perform a three-fold evaluation of generated images. Firstly, we distinguish between inherently visual stereotypes, which can be represented in an image, and ‘non-visual’ stereotypes. Subsequently, we identify visual stereotypes prevalent in the generated images from T2I models across 135 identity groups by conducting a large-scale annotation task. We will publicly release our dataset ‘ViSAGe’ which contains visual stereotypes and the annotated images featuring highlighted visual markers of these stereotypes. Additionally, we quantify the offensiveness of the generated images across different identity groups, and investigate the feasibility of automatically detecting stereotypes in images. Through an extensive case study, we also demonstrate that generated images of almost all identity groups exhibit a more visually ‘stereotypical’ appearance, even when Text-to-Image models are explicitly prompted with neutral or non-stereotypical attributes. This phenomenon is more prominent for identity groups from the global south. We hope that the presented approach and the dataset will provide valuable insights for understanding and addressing regional stereotypes in visual content generation.
@@
We evaluate the stereotypes generated in visual depictions of people, based on their identity associated with a geographical location. We acknowledge that this is a complex notion of identity and will have overlap with other facets of a person’s identity, such as race, which is not a focus of this work. We also emphasize that this work does not cover all possible visual depictions of stereotypes and thus, it’s methods or dataset should not be used as a standard to deem model generations free of stereotypes. This work is an initial step towards a systematic, global-scale evaluations of T2I models for stereotypical depictions, and we hope for future work to build on this to attain more coverage of potential stereotypical and harmful depictions in generations.
@@ Acknowledgements
We thank Utsav Prabhu, and Kathy Meier-Hellstern for their feedback and advice.